Tuesday, July 20, 2010

Problems and solutions

For some two or three weeks, we've been running into a problem with the experiment...it hasn't been giving us good data. We looked all over the place for answers...could it be a problem with the programming? No. The assembly itself? Maybe. After some time and bashing of ideas (mostly done by Drs Hart and Peterson) we came up with the idea that the support rings on either side of the potential assembly, made of an aluminum oxide, were charging up like a capacitor during the firing phase of our experiment. Since the aluminum has a decay constant of about three minutes, we waited for about that much time and found much better data for the dump of the plasma. The conclusion was to replace those rings with copper, a substance much less likely to charge up and ruin our plasma.

So we had new rings created. They were finished today and delivered to the plasma lab  at about two in the afternoon by Wes, the resident engineer at the physics department. Afterword, Dr Peterson, Chad and I worked on removing the central ring assembly, replacing the rings, then re-installing the assembly. We had to disconnect the wires leading to the potential assembly and haven't yet been able to reconnect them (that will be an adventure for tomorrow), but theoretically our data should be looking much better after we connect those lead wires and pump the system back down to the pressures we need (something like 10^-9 torr)

No comments:

Post a Comment